Mass 1 = 1.00kg

Mass 2 = 2.00 kg

Mass 1 initial velocity = 2.00 m/s

Mass 2 initial velocity = -1.0 m/s

Initial Total Velocity: ________??????

Final velocity of Mass 1 = -2.0 m/s

final velocity of mass 2 = 1.0 m/s

Final total velocity __________????
Answer #1

Solution) M1 = 1 kg

M2 = 2 kg

U1 = 2 m/s

U2 = - 1 m/s

Initial total velocity , U = ?

U = U1 + U2 = 2 - 1 = 1 m/s

U = 1 m/s

V1 = - 2 m/s

V2 = 1 m/s

Final total velocity , V = ?

V = V1 + V2 = - 2 + 1 = - 1 m/s

V = - 1 m/s